Please help me solve this expression:
1/3y + 4 ÷ 2 = 5, the value of y is 21.(a)True(b)False

Answer by rothauserc(4718)

1/3y + 4 ÷ 2 = 5
I assume your equation is
(1/3)y + 4 / 2 = 5
do all multiplication and division from left to right in the order they appear,
then addition and subtraction from left to right in the order they appear
(1/3)y + 2 = 5
(1/3)y = 3
multiply both sides of = by 3
y = 9
substitute in equation
(1/3)9 + 4 / 2 = 5
3 + 2 = 5
5 = 5
answer is (b)